OpenStudy (anonymous):

If x≠0 and x is inversely proportional to y, which of the following is directly proportional to 1/x^2? -1/y^2 1/y^2 1/y y y^2

OpenStudy (anonymous):

y2

OpenStudy (anonymous):

@vadevalor can u explain how u got that, bro?

OpenStudy (anonymous):

x = 1/y 1/x2 = 1/ (1/y)2=y2

OpenStudy (cwrw238):

yes thats right - i just got same result on paper

OpenStudy (cwrw238):

you substitute x = 1/y into 1/x^2 inversely proportion is y = k/x direst priportion is y = kx vadevalor used k = 1 in his proof which is ok